In a certain electronic factory, the ratio of the number of male to female workers is 2:3. If one hundred new female workers are hired, the number of female workers will increase to 68% of the total number of workers. Find the original number of workers in the factory.
Answer by venugopalramana(3286)

In a certain electronic factory, the ratio of the number of male to female workers is 2:3. If one hundred new female workers are hired, the number of female workers will increase to 68% of the total number of workers. Find the original number of workers in the factory.
ratio of the number of male to female workers is 2:3
so let males =2x
and females=3x
if 100 mnew females are taken then total females =3x+100
total workers =2x+3x+100=5x+100
% females=100*females/total=100*(3x+100)/(5x+100)=68 given.hence
100(3x+100)=68(5x+100)
300x+10000=340x+6800
340x-300x=10000-6800=3200
40x=3200..or..x=3200/40=80
hence original males are =2x=2*80=160
hence original females are =3x=3*80=240